हिंदी

If P(A) = 25, P(B) = 310 and P(A ∩ B) = 15, then P(A|B).P(B'|A') is equal to ______.

Advertisements
Advertisements

प्रश्न

If P(A) = `2/5`, P(B) = `3/10` and P(A ∩ B) = `1/5`, then P(A|B).P(B'|A') is equal to ______.

विकल्प

  • `5/6`

  • `5/7`

  • `25/42`

  • 1

MCQ
रिक्त स्थान भरें
Advertisements

उत्तर

If P(A) = `2/5`, P(B) = `3/10` and P(A ∩ B) = `1/5`, then P(A|B).P(B'|A') is equal to `25/42`.

Explanation:

Given that: P(A) = `2/5`, P(B) = `3/10` and P(A ∩ B) = `1/5`

P(A') = `1 - 2/5 = 3/5`

P(B') = `1 - 3/10 = 7/10`

And P(A' ∩ B') = 1 – P(A ∪ B)

= 1 – [P(A) + P(B) – P(A ∩ B)]

= `1 - [2/5 + 3/10 - 1/5]`

= `1 - [1/5 + 3/10]`

= `1 - 5/10`

= `1/2`

∴ `"P"("A'"/"B'") = ("P"("A'" ∩ "B'"))/("P"("B'"))`

= `(1/2)/(7/10)`

= `5/7`

And `"P"("B'"/"A'") = ("P"("A'" ∩ "B'"))/("P"("A'"))`

= `(1/2)/(3/5)`

= `5/6`

∴ `"P"("A'"/"B'")*"P"("B'"/"A'") = 5/7 xx 5/6`

= `25/42`

shaalaa.com
  क्या इस प्रश्न या उत्तर में कोई त्रुटि है?
अध्याय 13: Probability - Exercise [पृष्ठ २७९]

APPEARS IN

एनसीईआरटी एक्झांप्लर Mathematics Exemplar [English] Class 12
अध्याय 13 Probability
Exercise | Q 59 | पृष्ठ २७९

वीडियो ट्यूटोरियलVIEW ALL [2]

संबंधित प्रश्न

Assume that the chances of a patient having a heart attack is 40%. Assuming that a meditation and yoga course reduces the risk of heart attack by 30% and prescription of certain drug reduces its chance by 25%. At a time a patient can choose any one of the two options with equal probabilities. It is given that after going through one of the two options, the patient selected at random suffers a heart attack. Find the probability that the patient followed a course of meditation and yoga. Interpret the result and state which of the above stated methods is more beneficial for the patient.


40% students of a college reside in hostel and the remaining reside outside. At the end of the year, 50% of the hostelers got A grade while from outside students, only 30% got A grade in the examination. At the end of the year, a student of the college was chosen at random and was found to have gotten A grade. What is the probability that the selected student was a hosteler ?


If P(A) = 0.8, P(B) = 0.5 and P(B|A) = 0.4, find  P(A ∪ B)


Evaluate P(A ∪ B), if 2P(A) = P(B) = `5/13` and P(A | B) = `2/5`


Determine P(E|F).

A coin is tossed three times, where

E: at most two tails, F: at least one tail


Determine P(E|F).

Two coins are tossed once, where 

E: tail appears on one coin, F: one coin shows head


Determine P(E|F).

Mother, father and son line up at random for a family picture

E: son on one end, F: father in middle


Two balls are drawn at random with replacement from a box containing 10 black and 8 red balls. Find the probability that

  1. both balls are red.
  2. first ball is black and second is red.
  3. one of them is black and other is red.

A and B are two events such that P (A) ≠ 0. Find P (B|A), if A ∩ B = Φ.


A die is thrown again and again until three sixes are obtained. Find the probability of obtaining the third six in the sixth throw of the die.


In an examination, 30% of students have failed in subject I, 20% of the students have failed in subject II and 10% have failed in both subject I and subject II. A student is selected at random, what is the probability that the student has failed in at least one subject?


Two balls are drawn from an urn containing 5 green, 3 blue, and 7 yellow balls one by one without replacement. What is the probability that at least one ball is blue?


From a pack of well-shuffled cards, two cards are drawn at random. Find the probability that both the cards are diamonds when the first card drawn is replaced in the pack


Two cards are drawn one after the other from a pack of 52 cards without replacement. What is the probability that both the cards drawn are face cards?


Can two events be mutually exclusive and independent simultaneously?


If P(A) = 0.5, P(B) = 0.8 and P(B/A) = 0.8, find P(A/B) and P(A ∪ B)


One bag contains 5 white and 3 black balls. Another bag contains 4 white and 6 black balls. If one ball is drawn from each bag, find the probability that one white and one black


Two thirds of students in a class are boys and rest girls. It is known that the probability of a girl getting a first grade is 0.85 and that of boys is 0.70. Find the probability that a student chosen at random will get first grade marks.


Given P(A) = 0.4 and P(A ∪ B) = 0.7 Find P(B) if A and B are mutually exclusive


Given P(A) = 0.4 and P(A ∪ B) = 0.7 Find P(B) if A and B are independent events


A year is selected at random. What is the probability that it contains 53 Sundays


A year is selected at random. What is the probability that it is a leap year which contains 53 Sundays


Choose the correct alternative:

A, B, and C try to hit a target simultaneously but independently. Their respective probabilities of hitting the target are `3/4, 1/2, 5/8`. The probability that the target is hit by A or B but not by C is


In a multiple-choice question, there are three options out of which only one is correct. A person is guessing the answer at random. If there are 7 such questions, then the probability that he will get exactly 4 correct answers is ______ 


The total number of ways in which 5 balls of different colours can be distributed among 3 persons so that each person gets at least one ball is ______ 


Let A and B be two events. If P(A) = 0.2, P(B) = 0.4, P(A ∪ B) = 0.6, then P(A|B) is equal to ______.


If P(A ∩ B) = `7/10` and P(B) = `17/20`, then P(A|B) equals ______.


If P(A) = `3/10`, P(B) = `2/5` and P(A ∪ B) = `3/5`, then P(B|A) + P(A|B) equals ______.


Let A and B be two non-null events such that A ⊂ B. Then, which of the following statements is always correct?


For a biased dice, the probability for the different faces to turn up are

Face 1 2 3 4 5 6
P 0.10 0.32 0.21 0.15 0.05 0.17

The dice is tossed and it is told that either the face 1 or face 2 has shown up, then the probability that it is face 1, is ______.


Read the following passage:

Recent studies suggest the roughly 12% of the world population is left-handed.

Depending upon the parents, the chances of having a left-handed child are as follows:

A :  When both father and mother are left-handed:
Chances of left-handed child is 24%.
B :  When father is right-handed and mother is left-handed:
Chances of left-handed child is 22%.
C :  When father is left-handed and mother is right-handed:
Chances of left-handed child is 17%.
D :  When both father and mother are right-handed:
Chances of left-handed child is 9%.

Assuming that P(A) = P(B) = P(C) = P(D) = `1/4` and L denotes the event that child is left-handed.

Based on the above information, answer the following questions:

  1. Find `P(L/C)` (1)
  2. Find `P(overlineL/A)` (1)
  3. (a) Find `P(A/L)` (2)
    OR
    (b) Find the probability that a randomly selected child is left-handed given that exactly one of the parents is left-handed. (2)

Compute P(A|B), if P(B) = 0.5 and P (A ∩ B) = 0.32.


If A and B are any two events of S and \[P(F)\neq 0\], which formula is correct?


Share
Notifications

Englishहिंदीमराठी


      Forgot password?
Use app×